OpenStudy (annie96):

did I do this right. Simplify the expression –4(x – 7) + 3(x + 9). -4(x-7)+3(x+9) -4x-(-28)+3x+27 -4x+3x -1x-(-28)+27 27x+27=54x

OpenStudy (anonymous):

have you simplified it or do you need help with it

OpenStudy (annie96):

need help

OpenStudy (anonymous):

ok take the 27x to the right hand side. 27=54x-27x

OpenStudy (anonymous):

27=27x

OpenStudy (annie96):

okay

OpenStudy (anonymous):

then divide the x by 27

OpenStudy (annie96):

alright

OpenStudy (anonymous):

and 27 by 27 so 27x/27=27/27

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

now to check if the answer is correct substitute the x=1 into the original equation

OpenStudy (anonymous):

my bad i solved it instead of simplifying it sorry

OpenStudy (annie96):

okay

OpenStudy (anonymous):

give me 2 mins

OpenStudy (anonymous):

so we have 27x-27=54x

OpenStudy (anonymous):

we take the 27x and 27 to the right hand side

OpenStudy (anonymous):

so we have 54x-27x-27=0 27x-27=0 9(3x-3)=0 and that is the simplified version

OpenStudy (anonymous):

satisfied?